Determine whether the ordered pair is a solution of the given linear equation.
2x + 5y = -8; (-2,-2)

Answer by umnikmo(75)

lets substitute the given pair into the equation:
2*(-2)+5*(-2)=-4-10=-14, which is not equal to -8, so the pair is not a solution.